OpenStudy (anonymous):

Properties of Equality 8 = c/-4 + 6?

OpenStudy (anonymous):

\[8 = \frac{c}{-4} + 6\] \[8 -6 = \frac{c}{-4}\] \[2 = \frac{c}{-4}\] c = 2 x -4 c= -8

OpenStudy (anonymous):

._. dang...lol At least now I know how simple this was...Thank you :)